# LongVLM: Efficient Long Video Understanding via Large Language Models

### v2

@@ Abstract
Empowered by Large Language Models (LLMs), recent advancements in VideoLLMs have driven progress in various video understanding tasks. These models encode video representations through pooling or query aggregation over a vast number of visual tokens, making computational and memory costs affordable. Despite successfully providing an overall comprehension of video content, existing VideoLLMs still face challenges in achieving detailed understanding in videos due to overlooking local information in long-term videos. To tackle this challenge, we introduce LongVLM, a straightforward yet powerful VideoLLM for long video understanding, building upon the observation that long videos often consist of sequential key events, complex actions, and camera movements. Our approach proposes to decompose long videos into multiple short-term segments and encode local features for each local segment via a hierarchical token merging module. These features are concatenated in temporal order to maintain the storyline across sequential short-term segments. Additionally, we propose to integrate global semantics into each local feature to enhance context understanding. In this way, we encode video representations that incorporate both local and global information, enabling the LLM to generate comprehensive responses for long-term videos. Experimental results on the VideoChatGPT benchmark and zero-shot video question-answering datasets demonstrate the superior capabilities of our model over the previous state-of-the-art methods. Qualitative examples demonstrate that our model produces more precise responses for long videos understanding. Code will be available at https://github.com/ziplab/LongVLM.
@@ 1 Introduction
Figure 1: Previous approaches employ pooling operation or query aggregation for visual representations, focusing primarily on modeling global semantics and potentially overlooking detailed information in long-term videos. In contrast, our LongVLM proposes to aggregate local features for short-term segments and integrate global semantics into segment-level local feature sequence. Text highlighted in bold green denotes correct content, while text in red indicates errors. The video duration is 1 minute and 20 seconds.
@@
Large language models (LLMs) [36, 37, 45, 46, 43, 7] have revolutionized natural language understanding tasks and have demonstrated a remarkable capability to follow human instructions and intentions, emerging as a universal agent for general-purpose assistants. Drawing from the development of LLMs, Multi-modal Large Language Models (MLLMs) [30, 65, 8, 15] have driven the advance in vision-language learning by integrating visual encoders with LLMs and finetuning on language-image instruction-following data. However, developing Video-based Large Language Models (VideoLLMs) still poses a significant challenge due to the necessity of processing a large number of tokens for jointly modeling spatial-temporal dependencies across consecutive video frames. For instance, employing OpenAI CLIP-ViT-L/14 [40] as a visual encoder for a 100-frame video clip necessitates handling 25.6K visual tokens, leading to impractical computational costs with existing LLMs. To address this, recent approaches propose to extract video representation via precompression over visual tokens, utilizing pooling operation [35, 34] or query aggregation [24, 61, 42] over the video token sequence before feeding them into the LLM, as shown in Fig. 0(a). While these models showcase impressive capabilities in providing a meaningful understanding of video content, they still face challenges in achieving a significant advantage in fine-grained understanding of long-term videos. For example, as shown in Fig. 0(b), while all models recognize the overall environment (workshop), the object (bike), and the action (fixing), previous methods may fail to correctly identify details such as the color of the clothes (brown), or the specific component being fixed (bicycle chain).
@@
The main reason is that long-term videos typically involve numerous key actions, complex activities, and camera movements. Consequently, a long video can be divided into a sequence of short-term segments. For instance, in the example depicted in Fig. 0(b), various short-term actions occur, e.g. speaking, displaying spare components, grabbing the bicycle chain, along with the camera moving from the human to the bicycle wheel, and eventually focusing on the broken chain. Similarly, prior methods in video recognition task suggest to decompose complex activities into sequences of sub-activities [47, 19, 14]. These approaches treat the features of each short-range activity as the local information within the videos and emphasize the importance of reasoning over local features to develop a temporal-structural understanding [47, 14, 48, 49, 19, 60] within long-term videos for comprehending fine-grained information. From this perspective, existing VideoLLMs treat all visual tokens equally and aggregate them into compact representations through pooling operations [35, 34] and query aggregation [24, 61, 42]. While they successfully capture the global semantic context spanning the entire long-term videos, they often overlook preserving the local information for the short-term segments and the temporal structure of different short-term components, e.g., the order of events or sub-actions. However, exclusively modeling the temporal structure through the sequence of local features may still lead to inconsistent recognition across different segments and impede the overall understanding of the videos. To comprehend the content in long videos, the human visual system relies on a blend of local and global information [44]. Building on this insight, earlier approaches in video object detection [52, 54] suggest integrating global semantics into local localization descriptors, motivating us to include global semantic information into the sequence of local features for enriching the context understanding for each short-term segment.
@@
In this paper, we present LongVLM, a simple yet effective VideoLLM for efficient long video understanding, as illustrated in Fig. 2. We propose to extract video representations as sequences of short-term local features, and integrate global semantics into each short-term segment feature. Specifically, we begin by uniformly sampling a sequence of video frames from long-term videos and utilize a pretrained visual encoder, e.g., CLIP-ViT-L/14 [40], to extract visual features for each individual video frame. These frame-level features include the [CLS] tokens from a range of encoder layers and the patch features from the last second layer of the visual encoder. Then, we divide the sequence of patch features along the temporal dimension, resulting in a sequence of short-term segments. Each segment is considered as a local unit in the videos and includes patch features of the video frames within that segment. To reduce computational costs and obtain the compact features for each segment, a token merging module is employed to aggregate these patch features for the specific segment into a condensed set of tokens. In this way, we obtain the local features for each segment. We next concatenate these features sequentially to explicitly preserve the temporal order of the short-term segments in long-term videos. Moreover, we average the [CLS] tokens from each video frame along the temporal dimension to represent the global semantic information of the entire video. To integrate the global information, we sequentially arrange the averaged [CLS] tokens and the segment-level features, and then feed them into the LLM after passing through a projection layer. Benefiting from the causal attention mechanism in the LLM, we simultaneously achieve temporal structure modeling over the sequence of short-term segments and inject global semantics into the local features. Finally, the LLM generates responses based on the input sequence, which is composed of the obtained video representation and the designed system command with the specific user queries.
@@ 2.2 Video-based Large Language Models
Building upon pretrained image-language models, traditional approaches [33, 13, 11, 22, 50, 28, 6, 26] are proposed to address various video-language learning tasks.With the advent of LLMs, recent approaches are exploring the potential of developing Video-based Large Language Models (VideoLLMs) to unify various video-language understanding scenarios through human-video dialogue interactions. Existing VideoLLMs typically follow a common paradigm, which involves using a pretrained visual encoder to encode visual features, a projection layer to convert visual representations into the text latent space of LLMs, and a pretrained LLM for response generation. VideoChatGPT [35] and Valley [34] rely on pooling over visual tokens to obtain compact visual representations. VideoChat [24] utilizes pretrained video foundation models [25, 50] and Q-Former from BLIP-2 [23] to aggregate video representations. Video-LLaMA [61] proposes a Video Q-Former and an Audio Q-Former, enabling multiple modalities for video comprehension, while Video-ChatCaptioner [5] employs ChatGPT [36] to summarize video descriptions in multiple rounds of interactive question-and-answer conversation. Recently, MovieChat [42] proposes an effective memory management mechanism to enable LLMs to reason over hour-long videos. Multiple video-centric instruction datasets [34, 24, 35, 42] have also been proposed to finetune VideoLLMs for better video understanding capacity. Moreover, BT-Adapter [31] proposes a temporal adapter alongside the visual encoder for post-pretraining, while Video-Teller [29] highlights the importance of modality alignment in pretraining. Overall, these VideoLLMs rely on pooling and query aggregation on the whole long videos to extract visual representation for developing VideoLLMs, which lack of local information for fine-grained understanding in long videos. In contrast to existing VideoLLMs, we propose a simple yet effective framework that is feasible for aggregating both local and global information in long-term videos and preserves fine-grained content understanding.
@@ 2.3 Long-term Video Processing
Long-term video understanding poses several challenges due to the need to exploit complicated spatial-temporal dependencies while removing temporal redundancy over extended time duration. Previous studies propose efficient architectures [9, 20], temporal pooling/aggregation [12, 41, 64, 53], dynamic clip selection [21, 17, 16] to aggregate video representation while removing redundant information in videos. Other methods in video-language understanding tasks suggest to capture event temporality, causality, and dynamics in long-term videos by designing temporal alignment modules [18, 3]. Memory mechanism is also widely adopted in video dense prediction tasks [32, 52, 58, 52, 54, 63] to capture historical information and maintain temporal coherence, which results in more accurate and consistent prediction over time in long-term videos. In contrast to previous methods, we investigate the techniques for long-term video understanding in VideoLLMs. We propose to aggregate both local segment-level information and global semantic information, empowering MLLMs enhanced fine-grained understanding for long-term videos.
@@ 3 Method
In Sec. 3.1, we introduce the overall architecture and generation pipeline of the proposed LongVLM. In Sec. 3.2, we introduce the process of constructing local representation via short-term feature aggregation. In Sec. 3.3, we discuss the integration of both local segment-level feature and global semantic feature for the video representation.
@@ 3.1 Overall Architecture
Given an input video ${\mathcal{V}}\in{\mathbb{R}}^{T\times H\times W\times 3}$, we employ a visual encoder to extract frame-level features $\{{\bf X}^{t},{\bf P}^{t}\}_{t=1}^{T}$ for each video frame independently. Following previous methods [30, 35], we utilize the patch feature ${\bf P}^{t}\in{\mathbb{R}}^{N\times d}$ from the penultimate encoder layer for each video frame, where $N,d$ are the number of patch tokens and the channel dimension of the visual encoder, respectively. Additionally, we gather the [CLS] tokens ${\bf X}^{t}\in{\mathbb{R}}^{E\times d}$ from $E$ selected encoder layers for each individual video frame.
@@
We forward the concatenated global features and the sequence of local features into a projection layer ${\mathcal{F}}_{V}(\cdot;{\bf W})$ to obtain the projected visual feature, i.e., ${\bf H}_{V}=[\hat{{\bf G}},\hat{{\bf L}}]=[{\mathcal{F}}_{V}({\bf G};{\bf W}),{\mathcal{F}}_{V}({\bf L};{\bf W})]$. The projected visual feature ${\bf H}_{V}$ are concatenated with the tokenized system command and user queries ${\bf H}_{q}$, which are inputted into LLM for response generation ${\bf H}_{a}$, i.e., ${\bf H}_{a}={\mathcal{F}}_{LLM}([{\bf H}_{V},{\bf H}_{q}])$, where ${\mathcal{F}}_{LLM}(\cdot)$ denotes an LLM model.
@@ 3.2 Local Feature Aggregation
After obtaining the frame-level patch feature $\{{\bf P}^{t}\}_{t=1}^{T}$ from the last second layer of visual encoder, previous methods either apply factorized spatio-temporal pooling [35, 34], or utilize query aggregation [24, 61, 42] over all visual tokens, which may miss local information referring to the short-term events or actions. Nevertheless, videos have heavy spatio-temporal redundancy, which results in redundant computational costs by directly considering all the patch features as the local representation for each segment. Therefore, we propose to aggregate compact visual features within each short-term segment. Specifically, we collect the patch feature for the $s^{th}$ segment ${\bf V}^{s}=\{{\bf P}^{t}\}_{t=(s-1)K}^{sK}\in{\mathbb{R}}^{KN\times d}$ and apply a hierarchical token merging module to aggregate the local feature while reducing the number of visual tokens. Inspired by ToMe [1], we resort to the bipartite soft matching method and gradually merge the visual tokens for each short-term segment. At the $i^{th}$ step, we randomly partition the $R_{i}$ tokens into two non-overlap token sets ${\mathbb{P}}_{i}$ with $r_{i}$ tokens and ${\mathbb{Q}}_{i}$ with $R_{i}-r_{i}$ tokens, where initial $R_{0}=KN$. Then we calculate the similarity scores between the tokens in set ${\mathbb{P}}_{i}$ and ${\mathbb{Q}}_{i}$ based on the patch features. To obtain the similarity scores, each visual token is divided into $C$ heads along channel dimension, each with $\frac{d}{C}$ channels. The similarity score for each token pair is obtained by averaging the cosine similarity scores over all heads following
$a^{p_{i}q_{i}}=\frac{1}{C}[\sum_{c=1}^{C}{cos({\bf p}_{c}^{(p_{i})},{\bf p}_{c}^{(q_{i})})}],$
@@ 3.3 Global Semantics Integration
The local feature provides fine-grained information about different events or actions in the generation process of VideoLLMs, enhancing the detailed understanding capabilities of the LongVLM. However, the local feature for each segment may be insufficient for the model to reason the relationship between different segments and generate reasonable response over the entire videos. Therefore, we additionally introduce global semantic features to enrich the local feature sequence with contextual information. Specifically, we collect the [CLS] tokens from $E$ encoder layers from each video frame, i.e., $\{{\bf x}^{t}_{e}\}_{e=1,t=1}^{e=E,t=T}$, and then average the [CLS] tokens along temporal dimension, resulting in $\bar{{\bf X}}_{e}=\operatorname{AvgPool}(\{{\bf x}_{e}^{t}\}_{t=1}^{T})\in{\mathbb{R}}^{d},e\in[1,...,E]$. By default, $E$ can be the number of layers in the visual encoder. However, previous studies demonstrate different properties between intermediate features from the shallow layers and deeper layers in ViT models [27, 39], and showcase the deeper layers tend to aggregate global semantics. Thus, we concatenate the $E$-scale features along sequentially, resulting in the global semantic feature for the entire video, i.e., ${\bf G}=\{\bar{{\bf X}}_{e}\}_{e=1}^{E}\in{\mathbb{R}}^{E\times d}$.
@@
Remark. To address the risk of overlooking detailed understanding in long-term videos, we propose to divide long videos into multiple short-term segments and aggregate local spatial-temporal representation for each segment and preserving the temporal structure over the sequence of local feature vectors. Moreover, we enrich the local features with context information for better response generation by integrating global semantic information into short-term features. Different from previous approaches that rely solely on global semantics for long video understanding, we present a straightforward yet effective VideoLLM for achieving fine-grained understanding in long-term videos.

### v3

@@ Abstract
Empowered by Large Language Models (LLMs), recent advancements in Video-based LLMs (VideoLLMs) have driven progress in various video understanding tasks. These models encode video representations through pooling or query aggregation over a vast number of visual tokens, making computational and memory costs affordable. Despite successfully providing an overall comprehension of video content, existing VideoLLMs still face challenges in achieving detailed understanding due to overlooking local information in long-term videos. To tackle this challenge, we introduce LongVLM, a simple yet powerful VideoLLM for long video understanding, building upon the observation that long videos often consist of sequential key events, complex actions, and camera movements. Our approach proposes to decompose long videos into multiple short-term segments and encode local features for each segment via a hierarchical token merging module. These features are concatenated in temporal order to maintain the storyline across sequential short-term segments. Additionally, we propose to integrate global semantics into each local feature to enhance context understanding. In this way, we encode video representations that incorporate both local and global information, enabling the LLM to generate comprehensive responses for long-term videos. Experimental results on the VideoChatGPT benchmark and zero-shot video question-answering datasets demonstrate the superior capabilities of our model over the previous state-of-the-art methods. Qualitative examples show that our model produces more precise responses for long video understanding. Code is available at https://github.com/ziplab/LongVLM.
@@ 1 Introduction
Figure 1: (a) Comparison of model architectures. (b) Examples generated by different VideoLLMs. Text highlighted in bold green denotes correct content, while text in red indicates errors.
@@
Large language models (LLMs) [41, 1, 50, 51, 48, 8] have revolutionized natural language understanding tasks and have demonstrated a remarkable capability to follow human instructions and intentions, emerging as a universal agent for general-purpose assistants. Drawing from the development of LLMs, Multi-modal Large Language Models (MLLMs) [34, 70, 9, 16] have driven advancements in vision-language learning by integrating visual encoders with LLMs and finetuning on vision-language instruction pairs. However, developing Video-based Large Language Models (VideoLLMs) still poses a significant challenge due to the necessity of processing a large number of tokens for jointly modeling spatial-temporal dependencies across consecutive video frames. For instance, employing OpenAI CLIP-ViT-L/14 [44] as a visual encoder for a 100-frame video clip necessitates handling 25.6K visual tokens, leading to impractical computational costs with existing LLMs. To address this issue, recent approaches propose to extract video representation via precompression over visual tokens, utilizing pooling operation [39, 38] or query aggregation [28, 66, 47] over the video token sequence before feeding them into the LLM, as shown in Fig. 1(a). While these models showcase impressive capabilities in providing a meaningful understanding of video content, they still face challenges in achieving significant advantages in fine-grained understanding of long-term videos. For example, as shown in Fig. 1(b), while all models recognize the overall environment (workshop), the object (bike), and the action (fixing), previous methods may fail to correctly identify details such as the color of the clothes (brown), or the specific component being fixed (bicycle chain).
@@
The main reason is that long-term videos typically involve numerous key actions, complex activities, and camera movements. Consequently, a long video can be divided into a sequence of short-term segments. For instance, in the example depicted in Fig. 1(b), various short-term actions occur, e.g. speaking, displaying spare components, grabbing the bicycle chain, along with the camera moving from the human to the bicycle wheel, and eventually focusing on the broken chain. Similarly, prior methods in video recognition task suggest to decompose complex activities into sequences of sub-activities [52, 23, 15]. These approaches treat the features of each short-range activity as the local information within the videos and emphasize the importance of reasoning over local features to develop a temporal-structural understanding [52, 15, 53, 54, 23, 65] within long-term videos for comprehending fine-grained information. From this perspective, existing VideoLLMs treat all visual tokens equally and aggregate them into compact representations through pooling operations [39, 38] and query aggregation [28, 66]. While they successfully capture the global semantic context spanning the entire long-term videos, they often overlook preserving the local information for the short-term segments and the temporal structure of different short-term components, e.g., the order of events or sub-actions. However, exclusively modeling the temporal structure through the sequence of local features may still lead to inconsistent recognition across different segments and impede the overall understanding of the videos. To comprehend the content in long videos, the human visual system relies on a blend of local and global information [49]. Building on this insight, earlier approaches in video object detection [57, 59] suggest integrating global semantics into local localization descriptors, motivating us to include global semantic information into the sequence of local features for enriching the context understanding for each short-term segment.
@@
In this paper, we present LongVLM, a simple yet effective VideoLLM for efficient long video understanding, as illustrated in Fig. 2. We propose to extract video representations as sequences of short-term local features, and integrate global semantics into each short-term segment feature. Specifically, we begin by uniformly sampling a sequence of video frames from long-term videos and utilize a pretrained visual encoder, e.g., CLIP-ViT-L/14 [44], to extract visual features for each individual video frame. These frame-level features include the [CLS] tokens from a range of encoder layers and the patch features from the last second layer of the visual encoder. Then, we divide the sequence of patch features along the temporal dimension, resulting in multiple short-term segments. Each segment is considered as a local unit in the videos and includes patch features of the video frames within that segment. To reduce computational costs and obtain the compact features for each segment, a token merging module is employed to aggregate these patch features for the specific segment into a condensed set of tokens. In this way, we obtain the local features for each segment. We next concatenate these features sequentially to explicitly preserve the temporal order of the short-term segments in long-term videos. Moreover, we average the [CLS] tokens from each video frame along the temporal dimension to represent the global semantic information of the entire video. To integrate the global information, we prepend the averaged [CLS] tokens before the segment-level features, and then feed them into the LLM after passing through a projection layer. Benefiting from the causal attention mechanism in the LLM, we simultaneously achieve temporal structure modeling over the sequence of short-term segments and inject global semantics into the local features. Finally, the LLM generates responses based on the input sequence, which is composed of the obtained video representation and the designed system command with the specific user queries.
@@ 2.2 Video-based Large Language Models
Traditional video-language models [37, 14, 12, 26, 55, 32, 7, 30, 20] have advanced by using large-scale video-text pretraining followed by fine-tuning on specific downstream tasks. With the advent of LLMs, Video-based Large Language Models (VideoLLMs) explore various video-language understanding scenarios through human-video dialogue interactions. Existing VideoLLMs typically follow a common paradigm, which involves using a pretrained visual encoder to encode visual features, a projection layer to convert visual representations into the text latent space of LLMs, and a pretrained LLM for response generation. VideoChatGPT [39] and Valley [38] rely on pooling over visual tokens to obtain compact visual representations. VideoChat [28] utilizes pretrained video foundation models [29, 55] and Q-Former from BLIP-2 [27] to aggregate video representations. Video-LLaMA [66] proposes a Video Q-Former and an Audio Q-Former, enabling multiple modalities for video comprehension, while Video-ChatCaptioner [6] employs ChatGPT [41] to summarize video descriptions in multiple rounds of interactive question-and-answer conversation. Recently, MovieChat [47] proposes an effective memory management mechanism to enable LLMs to reason over hour-long videos. Multiple video-centric instruction datasets [38, 28, 39, 21] have also been proposed to finetune VideoLLMs for better video understanding capacity. Moreover, BT-Adapter [35] proposes a temporal adapter alongside the visual encoder for post-pretraining, while Video-Teller [33] highlights the importance of modality alignment in pretraining. Overall, these VideoLLMs rely on pooling and query aggregation on the whole long videos to extract visual representation, overlooking local information for fine-grained understanding in long videos. In contrast, we propose a simple yet effective framework that is feasible for aggregating both local and global information in long-term videos and preserves fine-grained content understanding.
@@ 2.3 Long-term Video Processing
Long-term video understanding poses several challenges due to the need to exploit complicated spatial-temporal dependencies while removing temporal redundancy over extended time duration. Previous studies propose efficient architectures [10, 24], temporal pooling/aggregation [13, 45, 69, 58], dynamic clip selection [25, 18, 17] to aggregate video representation while removing redundant information in videos. Other methods in video-language understanding tasks suggest to capture event temporality, causality, and dynamics in long-term videos by designing temporal alignment modules [22, 4]. Memory mechanism is also widely adopted in video dense prediction tasks [36, 57, 63, 59, 68] to capture historical information and maintain temporal coherence, which results in more accurate and consistent prediction over time in long-term videos. Differently, we propose to aggregate both local segment-level information and global semantic information, empowering MLLMs enhanced fine-grained understanding for long-term videos.
@@ 3 Method
In Sec. 3.1, we introduce the overall architecture and generation pipeline of the proposed LongVLM. In Sec. 3.2, we introduce the process of constructing local representation via short-term feature aggregation. In Sec. 3.3, we discuss the integration of both local segment-level feature and global semantic feature.
@@ 3.1 Overall Architecture
Given an input video ${\mathcal{V}}\in{\mathbb{R}}^{T\times H\times W\times 3}$, we employ a visual encoder to extract frame-level features $\{{\bf X}^{t},{\bf P}^{t}\}_{t=1}^{T}$ for each video frame independently. Following previous methods [34, 39], we utilize the patch feature ${\bf P}^{t}\in{\mathbb{R}}^{N\times d}$ from the second-to-last encoder layer, where $N,d$ are the number of patch tokens and the channel dimension of the visual encoder, respectively. Additionally, we gather the [CLS] tokens ${\bf X}^{t}\in{\mathbb{R}}^{E\times d}$ from $E$ selected encoder layers for each individual video frame.
@@
We forward the global features and the sequence of local features into a linear layer to obtain the projected visual features. The projected visual features are concatenated with the tokenized system command and user queries, which are inputted into LLM for response generation.
@@ 3.2 Local Feature Aggregation
After obtaining the frame-level patch feature $\{{\bf P}^{t}\}_{t=1}^{T}$ from the last second layer of visual encoder, previous methods either apply factorized spatio-temporal pooling [39, 38], or utilize query aggregation [28, 66] over all visual tokens, which may miss local information referring to the short-term events or actions. Nevertheless, videos have heavy spatio-temporal redundancy, which results in redundant computational costs by directly considering all the patch features as the local representation for each segment. Therefore, we propose to aggregate compact visual features within each short-term segment. Specifically, we collect the patch feature for the $s^{th}$ segment ${\bf V}^{s}=\{{\bf P}^{t}\}_{t=(s-1)K}^{sK}\in{\mathbb{R}}^{KN\times d}$ and apply a hierarchical token merging module to aggregate the local feature while reducing the number of visual tokens. Inspired by ToMe [2], we resort to the bipartite soft matching method and gradually merge the visual tokens for each short-term segment. At the $i^{th}$ step, we randomly partition the $R_{i}$ tokens into two non-overlap token sets ${\mathbb{P}}_{i}$ with $r_{i}$ tokens and ${\mathbb{Q}}_{i}$ with $R_{i}-r_{i}$ tokens, where initial $R_{0}=KN$. Then we calculate the similarity scores between the tokens in set ${\mathbb{P}}_{i}$ and ${\mathbb{Q}}_{i}$ based on the patch features. To obtain the similarity scores, each visual token is divided into $C$ heads along channel dimension, each with $\frac{d}{C}$ channels. The similarity score for each token pair is obtained by averaging the cosine similarity scores over all heads following Eq. 1:
$a^{p_{i}q_{i}}=\frac{1}{C}[\sum_{c=1}^{C}{cos({\bf p}_{c}^{(p_{i})},{\bf p}_{c}^{(q_{i})})}],$ (1)
@@ 3.3 Global Semantics Integration
The local features provide fine-grained information about different events or actions in the generation process of VideoLLMs, enhancing the detailed understanding capabilities of the LongVLM. However, the local features for each segment may be insufficient for the model to reason the relationship between different segments and generate reasonable response over the entire videos. Therefore, we additionally introduce global semantic features to enrich the local features with contextual information. Specifically, we collect the [CLS] tokens of each video frame from $E$ encoder layers, i.e., $\{{\bf x}^{t}_{e}\}_{e=1,t=1}^{e=E,t=T}$, and then average the [CLS] tokens along temporal dimension, resulting in $\bar{{\bf X}}_{e}=\operatorname{AvgPool}(\{{\bf x}_{e}^{t}\}_{t=1}^{T})\in{\mathbb{R}}^{d},e\in[1,...,E]$. By default, $E$ can be the number of layers in the visual encoder. However, previous studies demonstrate different properties between intermediate features from the shallow layers and deeper layers in ViT models [31, 43], and showcase the deeper layers tend to aggregate global semantics. Thus, we concatenate the $E$-scale features along sequentially, resulting in the global semantic feature for the entire video, i.e., ${\bf G}=\{\bar{{\bf X}}_{e}\}_{e=1}^{E}\in{\mathbb{R}}^{E\times d}$.
@@
Remark. To address the risk of overlooking detailed understanding in long-term videos, we propose to divide long videos into multiple short-term segments and aggregate local spatial-temporal representation for each segment and preserving the temporal structure over the sequence of local feature vectors. Moreover, we enrich the local features with context information for better response generation by integrating global semantic information into short-term features.